OpenStudy (anonymous):

INTEGRATE: e^x/(1+e^2x)

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ e ^{x} }{ 1+e ^{2x}}\]

OpenStudy (anonymous):

dude...that is a hard one:)

OpenStudy (loser66):

I don't have time, just give you the guidance let u =e^x dv = (1/(1+e^(2x)dx du = e^xdx v = 2 ln(1+e^(2x) after replacing to solve it by part, you will get a term back to original one. Don't worry, just put everything on and move the like terms to one side, the other side is the answer. Have to go. I am sorry for not being fully helped

OpenStudy (anonymous):

ahhhh! integration by parts...how quickly I forget...

OpenStudy (anonymous):

I would have never got v

OpenStudy (anonymous):

\[\int\limits \frac{ e^x }{ 1+u^2 } \frac{ du }{ e^x }\] leaves you with \[\int\limits \frac{ 1 }{ 1+u^2 }du\]

OpenStudy (anonymous):

and at that point in my calculus studies we were allowed to use a table...